OpenStudy (anonymous):

Evaluate the expression for x=-3 xsquared+4x+3 A. -18 B.-16 C.0 D.6

OpenStudy (anonymous):

-3^2=9 4(-3)=-12 +3 9-12+3=-3+3=0

OpenStudy (anonymous):

Thank you